Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Probability

Expert replies
by pakaskwa » Thu Apr 02, 2009 12:10 am
A certain roller coaster has 3 cars, and a passenger is equally likely to ride in any 1 of the 3 cars each time that passenger rides the roller coaster. If a certain passenger is to ride the roller coaster 3 times, what is the probability that the passenger will ride in each of the 3 cars?

A. 0
B. 1/9
C. 2/9
D. 1/3
E. 1

My answer is not even in the list. Can someone help me?
Join the discussion
Source: — Problem Solving |

by DanaJ » Thu Apr 02, 2009 12:21 am
First time he rides the roller coaster the probability of him picking a cart he hasn't ridden before is 1, since it's the first time he rides it.
Second time he rides the roller coaster the probability of him picking a cart he hasn't ridden before is 2/3, since he's already ridden one of the three carts.
Third time he rides the roller coaster the probability of him picking a cart he hasn't ridden before is 1/3, since he's already ridden two of the carts.

So my answer will be 1*1/3*2/3 =2/9.
Join the discussion

by pakaskwa » Thu Apr 02, 2009 12:55 am
Thanks Dana. It makes sense.

I had difficulty because I translated the problem: there are 3 balls in a bag, each in different color. Pick up any one ball from the bag without looking into it, then put it back. What is the probability of picking 3 times and each time it's a different color?

I think it's the same as the original question (taking different car each time for 3 times). So my calculation is 1x1/3x1/3=1/9. What is the problem in my calculation or translation?
Join the discussion